OpenStudy (anonymous):

How many of the 1000 smallest positive integers are congruent to 1 modulo 9?

OpenStudy (zzr0ck3r):

1,10,19,28, 37, 46, 55.....

OpenStudy (zzr0ck3r):

I guess this does not help solve how many...

ganeshie8 (ganeshie8):

\[1\le 9k+1\le 1000\] where \(k\) is a nonnegative integer

OpenStudy (zzr0ck3r):

that should do it lol

OpenStudy (usukidoll):

you need a positive integer that can produce a remainder 1 in modulo 9.. err? try chinese remainder theorem???

OpenStudy (anonymous):

wth is that

OpenStudy (usukidoll):

ah no wait... it's one of those guessing type things.. try 10 is congruent to 1 mod 9

OpenStudy (usukidoll):

are you in number theory?

OpenStudy (anonymous):

yeah

OpenStudy (usukidoll):

ooh I just passed that course.. :)

ganeshie8 (ganeshie8):

hey if you see zzr's first reply, you just need to find the number of terms in that arithmetic progression

OpenStudy (usukidoll):

if we're in mod 9 we can't use 1 through 8... and 9 is 0 in mod 9

OpenStudy (anonymous):

so are u saying that i should do 1+9(n-1)<1000 and guess for n

OpenStudy (usukidoll):

I had a similar problem on my final exam only it was what is least positive integer for 5x = 97 mod 84. grrrr!

OpenStudy (usukidoll):

though it looks like your question needs more than 1 answer... so 10 is only part of it x_X

OpenStudy (zzr0ck3r):

Do you not understand @ganeshie8 solution? he solved it without anything fancy...

OpenStudy (usukidoll):

ahhh do we let k = 1,2,3,4.......

OpenStudy (anonymous):

from 1 to 9 there is 1 number so in 999 there is 111 numbers :D

OpenStudy (anonymous):

no 112

OpenStudy (anonymous):

if u use th equation 1+9(n-1) we see that 112 would be it not 111

OpenStudy (zzr0ck3r):

well \(1\le 9k+1\le 1000\iff 0\le k\le 100\) 100-0= 100

OpenStudy (anonymous):

1000 mod 9=1 so add it to 111

OpenStudy (anonymous):

yeah so 112

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yeah its 112

OpenStudy (anonymous):

does anyone need this problem or can i close it now?

OpenStudy (usukidoll):

you can close it :)

ganeshie8 (ganeshie8):

@UsukiDoll do you solve ur question by using eiclid gcd algorithm ?

ganeshie8 (ganeshie8):

*euclid

ganeshie8 (ganeshie8):

5x = 97 mod 84 guess there are several ways to solve this, i think euclid gcd algorithm is the straightforward one..

OpenStudy (usukidoll):

yeah I did... I got... -17 and 1 when I did the backwards Euclidean Algorithim I did remember using gcd.

OpenStudy (anonymous):

also fyi ganeshi8 ur way was correct, using that inequality statement

OpenStudy (usukidoll):

I did forwards and backwards EA

OpenStudy (usukidoll):

I guess I might have passed the final to get a B in the number theory course ^_^

ganeshie8 (ganeshie8):

nice :) i was just thinking about how chinese remaidner thm applies here.. thats all

OpenStudy (anonymous):

what is the chinese remainder theorm?

ganeshie8 (ganeshie8):

nvm i see it !

OpenStudy (usukidoll):

since 9 = 3 x 3 I figured x congruent to 1 mod 3 twice but that would take forever.

OpenStudy (anonymous):

chine's dont work here :O its an other problem

ganeshie8 (ganeshie8):

Makes sense ! 5x = 97 mod 84 we can solve below simpler system using chinese remainder thm 5x = 97 mod 3 5x = 97 mod 4 5x = 97 mod 7

OpenStudy (usukidoll):

well I'm done with school for now so my brain is in mmorpg land

OpenStudy (usukidoll):

UGH! WHAT THE Fffff :(. now I feel like ssssssssssss........thanks really. :'(

OpenStudy (usukidoll):

because I know what Chinese Remainder theorem is... but since I wasn't given a system of congruences, it would've been an incongruent solution type thing.... :'(

ganeshie8 (ganeshie8):

:) chinese remainder thm works as the numbers are coprime (3, 4, 7) 84 = 3*4*7

OpenStudy (usukidoll):

-______________________-!

OpenStudy (anonymous):

wth is the chinese remainder theorm ;(

OpenStudy (usukidoll):

it's a super long computation that requires 2 or more congruence systems.

ganeshie8 (ganeshie8):

hey chinese remainder thm is a method to solve system of congriences

OpenStudy (usukidoll):

you know that pos problem was the first question on my final >:O/// sooo gawd the fact that I could've used CRT.. damn it

OpenStudy (anonymous):

how old are u guys?

OpenStudy (usukidoll):

26

ganeshie8 (ganeshie8):

chinese remainder thm allows you to solve a system of linear congriences : \[a_1x\equiv b_1\pmod{c_1} \] \[a_2x\equiv b_2\pmod{c_2} \] \[a_3x\equiv b_3\pmod{c_3} \] \[\cdots \]

OpenStudy (anonymous):

oh, I'm 12

ganeshie8 (ganeshie8):

\[12\equiv 26\pmod{14}\]

OpenStudy (usukidoll):

O_X! what's a 12 year old doing learning number theory

OpenStudy (anonymous):

#Asian

ganeshie8 (ganeshie8):

they have a weird course by name "applied math" where they teach modular arithmetic, number systems, statistics, probability and evverything else in math in just 6 months

OpenStudy (anonymous):

@UsukiDoll what college did u go to.

OpenStudy (anonymous):

@ganeshie8 but some cities allow children to skip grades or go to diff places for a period (as in class) to learn higher math.

ganeshie8 (ganeshie8):

ikr :) I think you will understand chinese remainder thm easily and it is a very interesting thm look it up in google when free

OpenStudy (anonymous):

anyway ganeshie do u live in the northern or southern part of india. Just curious

ganeshie8 (ganeshie8):

see if you can try these hard problems on your own first (they are not hard once you know chinese remainder thm, then you can do them in less than 10 seconds! )

ganeshie8 (ganeshie8):

i live in bangalore and it is south india.. you live in india too ?

OpenStudy (anonymous):

119 and no I live in America but I am planning to go to India this year to help out in the slums.

OpenStudy (usukidoll):

OMG THAT pirate problem! That was in my book

OpenStudy (anonymous):

wth a pirate's problem?

OpenStudy (usukidoll):

omfg read what ganeshie's attachment is displaying

OpenStudy (usukidoll):

p.s. for the first problem I ate the eggs.

OpenStudy (anonymous):

lawl and second problem is 301

ganeshie8 (ganeshie8):

119 is answer for #8 ? its a classic problem lol i think all those probloems can be worked without chines remainder theorem.. but takes lot of time and donkey work!

ganeshie8 (ganeshie8):

let me check the answer key, one sec..

OpenStudy (usukidoll):

school is over.. christmas break for me......... so my solving ain't here XD

OpenStudy (anonymous):

for the third one the pirate's problem what should i do after this

OpenStudy (usukidoll):

you beat up the pirates and say give me coins

OpenStudy (anonymous):

x%17=R3 x%16=R10 x%15=R0

ganeshie8 (ganeshie8):

looks you nailed them !!! http://gyazo.com/b6f2ee82da6b39953bf3f95f2cd475af

OpenStudy (usukidoll):

wtf... hey Who is heck is DHL global and what is it doing with my Mathematical Biology supplemental book?

OpenStudy (anonymous):

wait can u explain the third one

OpenStudy (anonymous):

i know n(the number)=15x

ganeshie8 (ganeshie8):

@UsukiDoll you want to try explaining him as your memory is fresh il need to review..

OpenStudy (anonymous):

and all that other congruence statement

OpenStudy (anonymous):

but what to do after that

OpenStudy (usukidoll):

awww sweettttt my other has 3 day priority mail... I'm sorry what? that problem was from last year... wasurimapelleta gomennasai ^_^

OpenStudy (anonymous):

wait i got it anyway

OpenStudy (anonymous):

so u make all the congruence statements substituted inside, which causes u with 1 final equation and thats how u do it

ganeshie8 (ganeshie8):

yes this is a great start x%17=R3 x%16=R10 x%15=R0

OpenStudy (anonymous):

-150 +4080s.

ganeshie8 (ganeshie8):

we get this system : x = 17a + 3 x = 16b + 10 x = 15c

OpenStudy (anonymous):

and thats the final equation however u subsitute s for 1 since thats the smallest number of coins and so forth

OpenStudy (anonymous):

anyway so how's India

ganeshie8 (ganeshie8):

idk how you got -150 +4080s. im getting something else.. wil catchup after lunch... :)

ganeshie8 (ganeshie8):

you should get 3930 for pirates problem..

OpenStudy (dan815):

hey

OpenStudy (dan815):

what does 1 modulo 9 mean

OpenStudy (dan815):

is it remainedder 1 when u divide by 9

OpenStudy (usukidoll):

mod 9 is = 0,1,2,3,4,5,6,7,8 1 is the remainder yes... so we can't use 1-8 because those numbers are too small for mod 9... the #9 is 0 in mod 9 and the #10 is 1 in mod 9.

OpenStudy (dan815):

ohh okay!!